OpenStudy (anonymous):

verify sint-1/cost=cost-cott/cost+cott

OpenStudy (anonymous):

\[\frac{ \cos t-\cot t }{ \cos t+\cot t }=\frac{ \cos t-\frac{ \cos t }{ \sin t } }{ \cos t+\frac{ \cos t }{ \sin t } }\] \[=\frac{ \cos t \sin t-\cos t }{\cos t \sin t+\cos t }=\frac{ \cos t \left( \sin t-1 \right) }{ \cos t \left( \sin t+1 \right) }\] \[=\frac{ \sin t-1 }{\sin t+1 }\] check the statement. is it \[\frac{ \sin t-1 }{ \cos t }=\frac{ \cos t-\cot t }{\cos t+\cot t }\] ?

OpenStudy (anonymous):

Please carefully re-write the equation.